OpenStudy (anonymous):

Find all solutions for x: log2x+log2(x-6)=4

OpenStudy (anonymous):

Okay, we first have to merge our log expressions together: \[\log _{2} x+\log _{2}(x-6)=4\]\[\log _{2} x(x-6)=4\]\[\log _{2}\left( x ^{2}-6x \right)=4\]Now, can you take it from here? @kat50036

OpenStudy (anonymous):

where did the 2nd \[\log_{2} \] go

OpenStudy (anonymous):

Remember the log property we spoke of earlier. If two log expressions are added, then their interiors can be multiplied under a single log. The second log expression was simply merged with the first.

OpenStudy (anonymous):

\[\log_{2} x ^{2}-6\log_{x} =4\]

OpenStudy (anonymous):

Thats not right is it

OpenStudy (anonymous):

^_^ that's an entirely different equation.

OpenStudy (anonymous):

ugg

OpenStudy (anonymous):

Thanks for your help Cardinal_Carlos

OpenStudy (anonymous):

That's okay. We can start from where we left off: \[\log _{2}(x ^{2}−6x)=4\] From here, we simply use the logarithmic property and get: \[x ^{2} - 6x = 2^{4}\]which simplifies to: \[x ^{2} - 6x - 16 = 0\] Now, we just factor out from here: \[\left( x - 8 \right)\left( x + 2 \right) = 0\] Our solutions are:\[x = -2 ~~~and ~~~x = 8\]

OpenStudy (anonymous):

You're very welcome @kat50036 ^_^